# Computing Floquet quasienergies in finite and extended systems: Role of electromagnetic and quantum-geometric gauges

### v1

@@ Abstract
We present an approach to compute the Floquet quasienergy spectrum of time-periodic systems. The method allows to characterize the light-matter interaction in finite and extended structures by carefully addressing the resolution of the position operator. In periodic systems we discuss the role of the quantum-geometric gauge freedom of Bloch states and employ a Wannier-based scheme to compute the required matrix elements. As a consequence, the method is accurate and applicable to a broad range of systems, from atoms and molecules to cold atomic gases and materials described by density functional theory, as well as model systems. We demonstrate the applicability of the approach by studying two cases: a particle trapped in a one-dimensional box and the semiconducting material BC2N. We employ the first example to provide a numerical proof of the invariance of the Floquet quasienergy spectrum with respect to the choice of electromagnetic gauge. The analysis of BC2N then serves to illustrate the physical effects described by the quasienergies, such as multiphoton resonances, and their expected range of occurrence in real materials in terms of external electric field and frequency of the drive pulse.
@@ I Introduction
The Floquet theory [1, 2] offers a versatile theoretical tool for studying the time-evolution of periodically driven systems. With its aid, optical phenomena like multiphoton resonances [3, 4] or dynamic localization [5, 6, 7, 8] can be naturally described. Additionally, the Floquet formalism is also useful in the characterization of nontrivial phases [9, 10, 11], such as the fractional Chern insulator [12], or the geometric interpretation of nonlinear optical responses [13].
Over the past two decades, the Floquet theory has been intensively employed to elucidate the behavior of cold atoms, where periodically modulated optical lattices simulate the effects of time-dependent forces [14, 8, 15, 16]. In this field, Floquet theory proves instrumental in understanding the creation of synthetic electromagnetic fields [17, 18], the interaction-driven transition from bosonic superfluids to Mott insulator phases [19, 20] or the realization of dynamical structures such as Majorana fermions [21], among other effects.
@@
Despite progress on multiple theoretical aspects of the Floquet formalism, the growing complexity of the target systems, from cold atoms to condensed matter, requires the development and deployment of improved numerical tools as well. A parameter-free approach, capable of accounting for real-world systems under complex driving forces is necessary for an accurate description of the dynamics accessed in experiments. A few recent works have taken steps in this direction by incorporating concepts of the Floquet formalism into first-principles schemes commonly employed in density functional theory (DFT) [23, 33].
In this manuscript we provide a general numerical approach to compute two central objects in the Floquet formalism, namely the effective Floquet Hamiltonian and its eigenvalues, the so-called quasienergies. These describe relevant time-dependent phenomena such as the aforementioned multiphoton resonances and dynamical localization, and serve as a basis for multiple treatments within the Floquet theory [15, 34]. To this end, we develop a scheme based on Wannier functions that handles arbitrary driving forces and is applicable to a broad range of systems, from simple theoretical models to real-world materials described by DFT. This tool is therefore applicable to systems that are commonly studied in the cold atom and condensed matter communities.
The paper is organized as follows. First in Sec. II we review the basic properties of the Floquet formalism and formally prove the electromagnetic gauge covariance (invariance) of the effective Floquet Hamiltonian (quasienergies). We provide a numerical verification in Sec. III by considering the length and velocity gauges in the benchmark particle-in-a-box system. In Sec. IV we generalize the formalism to deal with extended systems, where we pay attention to the intricacies related to the position operator and the quantum-geometric gauge freedom [35]. We exemplify the use of our approach in the semiconducting material BC2N in Sec. V, and discuss relevant features of the quasienergy spectrum, such as avoided crossings between energy levels. The most technical parts of the work are described in the appendices.
II Time-evolution operator, Floquet Hamiltonian and Electromagnetic Gauge transformations
@@ II.1 Preliminaries
If we now apply a time-dependent electromagnetic field, the system is described by the minimal coupling Hamiltonian [36]
@@
The dynamic evolution of the wave function from an initial time $t_{0}$ to a final time $t$ can be expressed making use of the time-evolution operator $\hat{U}(t,t_{0})$:
|$\ket{\psi(t)}=\hat{U}(t,t_{0})\ket{\psi(t_{0})}.$ (4)
This quantity can be formally defined using the time-dependent Hamiltonian in Eq. (2) as [37]
@@
$\hat{U}(t,t_{0})=\hat{U}(t,t^{\prime})\hat{U}(t^{\prime},t_{0}),\quad\forall t^{\prime}\in(-\infty,\infty).$ (6)
@@ II.2 The Floquet picture
Let us now consider the time-dependent Hamiltonian of Eq. (2) satisfying a time-periodic constraint of the form
@@
where $T$ defines the period and $\omega$ its associated frequency. Any instant in time can be decomposed as $t=nT+t_{r},\;n\in\mathbb{Z}$ with $t_{r}\in[t_{0},t_{0}+T]$. Then, making use of the semigroup property in Eq. (6), one can rewrite the time-evolution operator as
$\hat{U}(t,t_{0})=\hat{U}(t_{r},t_{0})\left[\hat{U}(t_{0}+T,t_{0})\right]^{n}.$ (8)
The quantity inside brackets $\hat{U}(t_{0}+T,t_{0})$ is the one-cycle evolution operator; just as the translation operator by a lattice vector plays a prominent role in condensed-matter physics, so does $\hat{U}(t_{0}+T,t_{0})$ in the Floquet formalism [34]. Since it is a unitary operator, one can define the following Hermitian operator
$\hat{H}_{\text{F}}(t_{0})=i\frac{\hbar}{T}\text{log}\left[\hat{U}(t_{0}+T,t_{0})\right]$ (9)
known as the effective Floquet Hamiltonian. Its eigenvalues $\varepsilon_{n}$ are real and are commonly termed Floquet quasienergies. Because of the presence of the complex logarithm, these eigenvalues are not uniquely defined; if $\varepsilon_{n}$ is a valid eigenvalue of $\hat{H}_{\text{F}}(t_{0})$, then the family of eigenvalues
@@ II.3.1 Review of the general case
Considering the relation of Eq. (14) between the state vectors in different gauges, the time-evolution operator in Eq. (4) is seen to transform as
$\hat{U^{\prime}}(t,t_{0})=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t)\right]\hat{U}(t,t_{0})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{0})\right].$ (16)
@@ II.3.2 Gauge covariance
As an example, the mechanical momentum $\hat{\bm{\Pi}}=\hat{\bm{p}}-e\hat{\bm{A}}$ follows the transformation relation of Eq. (17), while the canonical momentum $\bm{\hat{p}}$ does not. In general, the Hamiltonian is not a gauge-covariant operator due to the presence of the last term in Eq. (15); only in the time-independent case does it transform covariantly, owing to $d\hat{\Lambda}/{dt}=0$, and the time-independent energy eigenvalues then are gauge-invariant. As for the time-evolution operator, in general it is not covariant due to the different time factors entering the exponentials in Eq. (16). However, this equation does ensure that its matrix elements are gauge-invariant, i.e., the probablity amplitude $\braket{\psi(t)}{\hat{U}(t,t_{0})}{\psi(t_{0})}=\braket{\psi^{\prime}(t)}{\hat{U}^{\prime}(t,t_{0})}{\psi^{\prime}(t_{0})}$ does not depend on the choice of gauge [38].
In the following, we show that in time-periodic systems the one-cycle evolution operator and, by extension, the Floquet Hamiltonian transform gauge-covariantly, which implies that the Floquet quasienergies are gauge-independent quantities.
@@ II.3.3 Time periodicity and the Floquet invariant
Let us reconsider the time-periodic condition of Eq. (7). For it to hold, both the vector- and scalar-gauge potentials must be time periodic,
@@
These conditions, in turn, impose a time-periodic restriction on the gauge transformation operator $\hat{\Lambda}(t)$ by virtue of Eq. (13):
$\begin{split}\hat{\Lambda}(t+T)=\hat{\Lambda}(t).\end{split}$ (19)
Equation (19) has some subtle consequences involving the physical conditions for the realization of Floquet compatible driving. We analyze this in App. A. This also has immediate consequences for the gauge transformation property of the one-cycle evolution operator $\hat{U}(t_{0}+T,t_{0})$ discussed in Sec. II.2. Setting $t=t_{0}+T$ in Eq. (16) we obtain
$\begin{split}&\hat{U}^{\prime}(t_{0}+T,t_{0})\\ &=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t_{0}+T)\right]\hat{U}(t_{0}+T,t_{0})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{0})\right]\\ &=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t_{0})\right]\hat{U}(t_{0}+T,t_{0})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{0})\right],\end{split}$ (20)
@@
$\begin{split}\hat{H}^{\prime}_{\text{F}}(t_{0})=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t_{0})\right]\hat{H}_{\text{F}}(t_{0})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{0})\right],\end{split}$ (21)
@@
We conclude this section with a brief remark on the gauge invariance of the quasienergies. The time-periodic condition of Eq. (19) is not the most general solution to the functional dependence problem of Eq. (18). The general solution is given by $\hat{\Lambda}(t+T)=\hat{\Lambda}(t)+\hat{\mathds{1}}\beta t$, where $\beta$ is a real constant with units of energy. In practice, $\beta$ acts as a rigid shift of the zero of the gauge potentials, which implies that only differences in quasienergies are gauge-invariant under a shift of the zero of energy. With this remark in place, given that $\beta$ is completely irrelevant for the dynamics and that the quasienergy spectrum is only uniquely defined within the quasienergy first Brillouin zone of Eq. (11), we will disregard it in the rest of the work and refer to quasienergies as gauge-invariant quantities.
@@ III.1 Electric dipole approximation: length and velocity gauges
$\begin{split}\hat{H}^{V}(t)=&\text{exp}\left[\frac{iq}{\hbar}\bm{A}(t)\cdot\hat{\bm{r}}\right]\hat{H}_{0}\;\text{exp}\left[\frac{-iq}{\hbar}\bm{A}(t)\cdot\hat{\bm{r}}\right].\end{split}$ (30)
In obtaining the above expression, the $d\hat{\Lambda}/{dt}$ term of Eq. (15) was cancelled by the position-operator term of Eq. (28).
@@ III.1.1 Remark on the velocity gauge
While correct in the ideal case, the Hamiltonian of Eq. (31) is only valid if one employs a complete set of bands and a Hamiltonian exactly expressible as Eq. (1) for the calculation of its matrix elements; but this is hardly never the case in practical calculations, where the set of basis states is usually truncated and therefore the form of $\hat{H}_{0}$ is no longer given by Eq. (1). A clarifying view on this issue has been provided recently in Refs. [40, 43]. The authors noted that, unlike in the length gauge, the form of the interaction term depends explicitly on $\hat{H}_{0}$ in the velocity gauge. As a consequence, a truncation of $\hat{H}_{0}$ implies a different form of the perturbation when using the expression in Eq. (31), leading to results that are incorrect. The correct expression for the Hamiltonian in the velocity gauge, even for a truncated set of states, is provided by Eq. (30). In App. B, we further analyse this point in the context of the breakdown of the commutator identity $[\hat{r}_{i},\hat{p}_{j}]=i\hbar\delta_{ij}\hat{\mathds{1}}$.
III.2 Quasi-energy dispersion
@@ III.2.1 Matrix elements and numerical details
with $(\hat{W})_{nm}=\left(\text{exp}\left[\frac{iq}{\hbar}\bm{A}(t)\cdot\hat{\bm{r}}\right]\right)_{nm}$. Finally, the matrix elements of the velocity-gauge Hamiltonian in the form of Eq. (31) are given by
$\begin{split}H^{V_{2}}_{nm}(t)=E_{n}\delta_{nm}-\frac{q}{m}\bm{A}(t)\cdot\bm{p}_{nm}+\dfrac{q^{2}A^{2}(t)}{2m}\delta_{nm},\end{split}$ (35)
@@
In all cases, we calculate the quasienergy spectrum using a discretized expression of the one-cycle evolution operator; the technical details regarding the numerical scheme can be found in App. C.

### v2

@@ Abstract
We present an approach to compute the Floquet quasienergy spectrum of time periodic systems. The method allows to characterize the light-matter interaction in finite and extended structures by carefully addressing the resolution of the position operator. In periodic systems we discuss the role of the quantum-geometric gauge freedom of Bloch states and employ a Wannier based scheme to compute the required matrix elements. As a consequence, the method is accurate and applicable to a broad range of systems, from atoms and molecules to cold atomic gases and materials described by density functional theory, as well as model systems. We demonstrate the applicability of the approach by studying two cases: a particle trapped in a one dimensional box and the semiconducting material BC2N. We employ the first example to provide a numerical proof of the invariance of the Floquet quasienergy spectrum with respect to the choice of electromagnetic gauge. The analysis of BC2N then serves to illustrate the physical effects described by the quasienergies, such as multiphoton resonances, and their expected range of occurrence in real materials in terms of external electric field and frequency of the drive pulse.
@@ I Introduction
The Floquet theory [1, 2] offers a versatile theoretical tool for studying the time evolution of periodically driven systems. With its aid, optical phenomena like multiphoton resonances [3, 4] or dynamic localization [5, 6, 7, 8] can be naturally described. Additionally, the Floquet formalism is also useful in the characterization of nontrivial phases [9, 10, 11], such as the fractional Chern insulator [12], or the geometric interpretation of nonlinear optical responses [13].
Over the past two decades, the Floquet theory has been intensively employed to elucidate the behavior of cold atoms, where periodically modulated optical lattices simulate the effects of time dependent forces [14, 8, 15, 16]. In this field, Floquet theory proves instrumental in understanding the creation of synthetic electromagnetic fields [17, 18], the interaction-driven transition from bosonic superfluids to Mott insulator phases [19, 20] or the realization of dynamical structures such as Majorana fermions [21], among other effects.
@@
Despite progress on multiple theoretical aspects of the Floquet formalism, the growing complexity of the target systems, from cold atoms to condensed matter, requires the development and deployment of improved numerical tools as well. A parameter-free approach, capable of accounting for real-world systems under complex driving forces is necessary for an accurate description of the dynamics accessed in experiments. A few recent works have taken steps in this direction by incorporating concepts of the Floquet formalism into first principles schemes commonly employed in density functional theory (DFT) [23, 33].
In this manuscript we provide a general numerical approach to compute two central objects in the Floquet formalism, namely the effective Floquet Hamiltonian and its eigenvalues, the so-called quasienergies. These describe relevant time dependent phenomena such as the aforementioned multiphoton resonances and dynamical localization, and serve as a basis for multiple treatments within the Floquet theory [15, 34]. To this end, we develop a scheme based on Wannier functions that handles arbitrary driving forces and is applicable to a broad range of systems, from simple theoretical models to real-world materials described by DFT. This tool is therefore applicable to systems that are commonly studied in the cold atom and condensed matter communities.
The paper is organized as follows. First in Sec. II we review the basic properties of the Floquet formalism and formally prove the electromagnetic gauge-covariance (invariance) of the effective Floquet Hamiltonian (quasienergies). We provide a numerical verification in Sec. III by considering the length and velocity gauges in the benchmark particle in a box system. In Sec. IV we generalize the formalism to deal with extended systems, where we pay attention to the intricacies related to the position operator and the quantum-geometric gauge freedom [35]. We exemplify the use of our approach in the semiconducting material BC2N in Sec. V, and discuss relevant features of the quasienergy spectrum, such as avoided crossings between energy levels. The most technical parts of the work are described in the appendices.
II Time evolution operator, Floquet Hamiltonian and Electromagnetic Gauge transformations
@@ II.1 Preliminaries
If we now apply a time dependent electromagnetic field, the system is described by the minimal coupling Hamiltonian [36]
@@
The dynamic evolution of the wave function from an initial time $t_{\text{s}}$ to a final time $t$ can be expressed making use of the time evolution operator $\hat{U}(t,t_{\text{s}})$:
|$\ket{\psi(t)}=\hat{U}(t,t_{\text{s}})\ket{\psi(t_{\text{s}})}.$ (4)
This quantity can be formally defined using the time dependent Hamiltonian in Eq. (2) as [37]
@@
$\hat{U}(t,t_{\text{s}})=\hat{U}(t,t^{\prime})\hat{U}(t^{\prime},t_{\text{s}}),\quad\forall t^{\prime}\in(-\infty,\infty).$ (6)
@@ II.2 The Floquet picture
Let us now consider the time dependent Hamiltonian of Eq. (2) satisfying a time periodic constraint of the form
@@
where $T$ defines the period and $\omega$ its associated frequency. Any instant in time can be decomposed as $t=nT+t_{r},\;n\in\mathbb{Z}$ with $t_{r}\in[t_{\text{s}},t_{\text{s}}+T]$. Then, making use of the semigroup property in Eq. (6), one can rewrite the time evolution operator as
$\hat{U}(t,t_{\text{s}})=\hat{U}(t_{r},t_{\text{s}})\left[\hat{U}(t_{\text{s}}+T,t_{\text{s}})\right]^{n}.$ (8)
The quantity inside brackets $\hat{U}(t_{\text{s}}+T,t_{\text{s}})$ is the one-cycle evolution operator; just as the translation operator by a lattice vector plays a prominent role in condensed-matter physics, so does $\hat{U}(t_{\text{s}}+T,t_{\text{s}})$ in the Floquet formalism [34]. Since it is a unitary operator, one can define the following Hermitian operator
$\hat{H}_{\text{F}}(t_{\text{s}})=i\frac{\hbar}{T}\text{log}\left[\hat{U}(t_{\text{s}}+T,t_{\text{s}})\right]$ (9)
known as the effective Floquet Hamiltonian. Its eigenvalues $\varepsilon_{n}$ are real and are commonly termed Floquet quasienergies. Because of the presence of the complex logarithm, these eigenvalues are not uniquely defined; if $\varepsilon_{n}$ is a valid eigenvalue of $\hat{H}_{\text{F}}(t_{\text{s}})$, then the family of eigenvalues
@@ II.3.1 Review of the general case
Considering the relation of Eq. (14) between the state vectors in different gauges, the time evolution operator in Eq. (4) is seen to transform as
$\hat{U^{\prime}}(t,t_{\text{s}})=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t)\right]\hat{U}(t,t_{\text{s}})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{\text{s}})\right].$ (16)
@@ II.3.2 Gauge-covariance
As an example, the mechanical momentum $\hat{\bm{\Pi}}=\hat{\bm{p}}-e\hat{\bm{A}}$ follows the transformation relation of Eq. (17), while the canonical momentum $\bm{\hat{p}}$ does not. In general, the Hamiltonian is not a gauge-covariant operator due to the presence of the last term in Eq. (15); only in the time independent case does it transform covariantly, owing to $d\hat{\Lambda}/{dt}=0$, and the time independent energy eigenvalues then are gauge-invariant. As for the time evolution operator, in general it is not covariant due to the different time factors entering the exponentials in Eq. (16). However, this equation does ensure that its matrix elements are gauge-invariant, i.e., the probablity amplitude $\braket{\psi(t)}{\hat{U}(t,t_{\text{s}})}{\psi(t_{\text{s}})}=\braket{\psi^{\prime}(t)}{\hat{U}^{\prime}(t,t_{\text{s}})}{\psi^{\prime}(t_{\text{s}})}$ does not depend on the choice of gauge [38].
In the following, we show that in time periodic systems the one-cycle evolution operator and, by extension, the Floquet Hamiltonian transform gauge-covariantly, which implies that the Floquet quasienergies are gauge-independent quantities.
@@ II.3.3 Time periodicity and the Floquet invariant
Let us reconsider the time periodic condition of Eq. (7). For it to hold, both the vector- and scalar-gauge potentials must be time periodic,
@@
These conditions, in turn, impose a time periodic restriction on the gauge transformation operator $\hat{\Lambda}(t)$ by virtue of Eq. (13):
$\hat{\Lambda}(t+T)=\hat{\Lambda}(t).$ (19)
Equation (19) has some subtle consequences involving the physical conditions for the realization of Floquet compatible driving. We analyze this in Appendix A. This also has immediate consequences for the gauge transformation property of the one-cycle evolution operator $\hat{U}(t_{\text{s}}+T,t_{\text{s}})$ discussed in Sec. II.2. Setting $t=t_{\text{s}}+T$ in Eq. (16) we obtain
$\begin{split}&\hat{U}^{\prime}(t_{\text{s}}+T,t_{\text{s}})\\ &=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t_{\text{s}}+T)\right]\hat{U}(t_{\text{s}}+T,t_{\text{s}})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{\text{s}})\right]\\ &=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t_{\text{s}})\right]\hat{U}(t_{\text{s}}+T,t_{\text{s}})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{\text{s}})\right],\end{split}$ (20)
@@
$\hat{H}^{\prime}_{\text{F}}(t_{\text{s}})=\text{exp}\left[\frac{iq}{\hbar}\hat{\Lambda}(t_{\text{s}})\right]\hat{H}_{\text{F}}(t_{\text{s}})\;\text{exp}\left[\frac{-iq}{\hbar}\hat{\Lambda}(t_{\text{s}})\right],$ (21)
@@
We conclude this section with a brief remark on the gauge-invariance of the quasienergies. The time periodic condition of Eq. (19) is not the most general solution to the functional dependence problem of Eq. (18). The general solution is given by $\hat{\Lambda}(t+T)=\hat{\Lambda}(t)+\hat{\mathds{1}}\beta t$, where $\beta$ is a real constant with units of energy. In practice, $\beta$ acts as a rigid shift of the zero of the gauge potentials, which implies that only differences in quasienergies are gauge-invariant under a shift of the zero of energy. With this remark in place, given that $\beta$ is completely irrelevant for the dynamics and that the quasienergy spectrum is only uniquely defined within the quasienergy first Brillouin zone of Eq. (11), we will disregard it in the rest of the work and refer to quasienergies as gauge-invariant quantities.
@@ III.1 Electric dipole approximation: length and velocity gauges
$\hat{H}^{V}(t)=\text{exp}\left[\frac{iq}{\hbar}\bm{A}(t)\cdot\hat{\bm{r}}\right]\hat{H}_{0}\;\text{exp}\left[\frac{-iq}{\hbar}\bm{A}(t)\cdot\hat{\bm{r}}\right].$ (30)
In obtaining the above expression, the $d\hat{\Lambda}/{dt}$ term of Eq. (15) was canceled by the position operator term of Eq. (28).
@@ III.1.1 Remark on the velocity gauge
While correct in the ideal case, the Hamiltonian of Eq. (31) is only valid if one employs a complete set of bands and a Hamiltonian exactly expressible as Eq. (1) for the calculation of its matrix elements; but this is hardly ever the case in practical calculations, where the set of basis states is usually truncated and therefore the form of $\hat{H}_{0}$ is no longer given by Eq. (1). A clarifying view on this issue has been provided recently in Refs. [40, 43]. The authors noted that, unlike in the length gauge, the form of the interaction term depends explicitly on $\hat{H}_{0}$ in the velocity gauge. As a consequence, a truncation of $\hat{H}_{0}$ implies a different form of the perturbation when using the expression in Eq. (31), leading to results that are incorrect. The correct expression for the Hamiltonian in the velocity gauge, even for a truncated set of states, is provided by Eq. (30). In Appendix B, we further analyze this point in the context of the breakdown of the commutator identity $[\hat{r}_{i},\hat{p}_{j}]=i\hbar\delta_{ij}\hat{\mathds{1}}$.
III.2 Quasienergy dispersion
@@ III.2.1 Matrix elements and numerical details
with $(\hat{W})_{nm}=\left(\text{exp}\left[\frac{iq}{\hbar}\bm{A}(t)\cdot\hat{\bm{r}}\right]\right)_{nm}$. Finally, the matrix elements of the velocity gauge Hamiltonian in the form of Eq. (31) are given by
$H^{V_{2}}_{nm}(t)=E_{n}\delta_{nm}-\frac{q}{m}\bm{A}(t)\cdot\bm{p}_{nm}+\dfrac{q^{2}A^{2}(t)}{2m}\delta_{nm},$ (35)
@@
In all cases, we calculate the quasienergy spectrum using a discretized expression of the one-cycle evolution operator; the technical details regarding the numerical scheme can be found in Appendix C.